OKust be shown in space proviaed. Remember units on all numbers I . Iron (III) chloride and hydrogen sulfide react. The balanced equation is: Given 3.45 mol of H2S and 2.55 mol FeCls, what is the limiting reactant? 2FeCl, + 3HaS → FeS, + 6HC1
Answer #1

2FeCl3 + 3H2S   \rightarrow Fe2S3 + 6HCl

According to the stoichiometric coefficient of the equation 2 mol of FeCl3eact completely with 3 mol of H2S.

Moles of H2S that will react with 1 mol of FeCl3 = 3/2 =1.5 mol

We are given 2.55 mol of FeCl3.

Moles of H2S that will react with 2.55 mol of FeCl3 = 2.55*(1.5)

= 3.825 mol

But we have only 3.45 mol of H2S , therefore H2S is the limiting reagent.